If force B does not change but force A is increased to 1200N this produces an acceleration of 0.5m/s squared Calculate the size of Force B. The mass of the car is 750kg


1
Expert's answer
2021-11-03T10:20:03-0400

"F=ma\to F_A+F_B=ma\to F_B=ma-F_A=750\cdot0.5-1200=-825\ (N)" . Answer